# Fluctuations in random field Ising models

### v1

@@ 3 Proof of main results
Proof of 2.9.
In this case we have
$\lVert\mathbf{A}_{n}\rVert_{\infty}=\frac{|\theta|}{2nh(1-h)}\max_{i\in[n]}d_{i}=\frac{|\theta|}{2\min(h,1-h)}<1,\quad\text{and}\quad\alpha_{n}=O(n^{-1}).$
Finally, for $t\in\{-1,1\}$ setting
$q_{i}=\begin{cases}\frac{1}{\sqrt{2nh}}&\mbox{if }1\leq i\leq nh\\ \frac{t}{\sqrt{2n(1-h)}}&\mbox{if }nh+1\leq i\leq n\end{cases},$
we have that $\Big{(}\mathbf{q},\frac{\theta t}{2\sqrt{h(1-h)}}\Big{)}$ forms an exact eigenvector-eigenvalue pair. The conclusion now follows immediately from 2.6; note that statement in the corollary rescales the eigenvectors by $\sqrt{2}$. ∎
@@
To prove 2.10, we require the following lemmas to check $\sqrt{n}\operatorname{\alpha_{\mathnormal{n}}}=o_{P}(1)$ as well as the MHT condition. Note that unlike the previous examples in the context of graphs, bounding the norm $\|\mathbf{A}_{n}\|_{4\to 4}$ by $\|\mathbf{A}_{n}\|_{\infty\to\infty}$ is suboptimal for spin glasses. We postpone the proofs of these lemmas to Section 6.
@@
Suppose $n\leq N$. Let $\mathbf{Z}$ be a $N\times n$ random matrix whose entries are independent sub-Gaussian random variables with mean zero and variance $1$, and sub-Gaussian norm bounded by a constant. For $r\in[2,\infty]$, the following bound holds for $\mathbf{A}_{n}:=\frac{1}{N}\operatorname{\mathrm{Off}}(\mathbf{Z}^{\top}\mathbf{Z})$:
$\|\mathbf{A}_{n}\|_{r\to r}=O_{P}\Big{(}\frac{n^{1-1/r}}{\sqrt{N}}\Big{)}.$ (4.9)
While not necessary for our analysis, we also have $\|\mathbf{A}_{n}\|_{r\to r}=O_{P}\Big{(}\frac{n^{1/r}}{\sqrt{N}}\Big{)}$ for $r\in[1,2)$.
@@
This result resembles Chevet’s inequality [11, 26], which provides similar $r\to r$ norm bounds for the case when $\mathbf{A}_{n}$ has i.i.d. entries. We are not aware of analogous results for covariance matrices, and this lemma may be of independent interest.
@@
Let $\mathbf{Z}$ be a $N\times n$ random matrix whose entries are independent sub-Gaussians with mean zero and variance $1$. Then, under the asymptotic scaling $N\gg n^{3/2}$, we have $\sqrt{n}\alpha_{n}=o_{P}(1).$
Proof of 2.10.
@@
Recall that we work under the asymptotic setting with $N\gg n^{3/2}$. By Lemma 4.2, we have $\sqrt{n}\operatorname{\alpha_{\mathnormal{n}}}=o_{P}(1)$. Also, by Lemma 4.1 with $r=4$, the MHT condition (2.2) holds. Finally, by Lemma 4.1 with $r=2$, we have $\lVert\mathbf{A}_{n}\rVert=O_{P}\left(\sqrt{\frac{n}{N}}\right),$ and so
$\sqrt{n\alpha_{n}}\lVert\mathbf{A}_{n}\mathbf{q}\rVert=o_{P}(n^{1/4})\lVert\mathbf{A}_{n}\mathbf{q}\rVert\leq o_{P}(n^{1/4})\lVert\mathbf{A}_{n}\rVert=o_{P}(1).$
@@ 5 Proof of main lemmas
(a). As in the proof of 2.4, we generate $\operatorname{\mathnormal{\bm{\sigma}}}^{\prime}$ as follows: let $I$ be a randomly sampled index from $\{1,2,\ldots,n\}$. Given $I=i$, replace $\sigma_{i}$ with an independently generated $\sigma^{\prime}_{i}$ drawn from the conditional distribution of $\sigma_{i}$ given $\sigma_{j},\ j\neq i$. Recall the definition of $\mathbf{b}^{(g)}$ from 3.1 and let
@@
$f(\operatorname{\mathnormal{\bm{\sigma}}}):=\operatorname{\mathbb{E}}[F(\operatorname{\mathnormal{\bm{\sigma}}},\operatorname{\mathnormal{\bm{\sigma}}}^{\prime})|\operatorname{\mathnormal{\bm{\sigma}}}]=\frac{1}{p}\sum_{\mathnormal{i}=1}^{\mathnormal{n}}\gamma_{i}(g(\sigma_{i})-b_{i}^{(g)}).$
Define $\Delta(\operatorname{\mathnormal{\bm{\sigma}}}):=\operatorname{\mathbb{E}}\left(|f(\operatorname{\mathnormal{\bm{\sigma}}})-f(\operatorname{\mathnormal{\bm{\sigma}}}^{\prime})||F(\operatorname{\mathnormal{\bm{\sigma}}},\operatorname{\mathnormal{\bm{\sigma}}}^{\prime})|\big{|}\operatorname{\mathnormal{\bm{\sigma}}}\right).$ By [8, Theorem 1.5(ii)], it suffices to show
@@
$f(\operatorname{\mathnormal{\bm{\sigma}}}^{\prime})=\frac{1}{p}\sum_{\mathnormal{i}=1}^{\mathnormal{n}}\gamma_{i}(g(\sigma_{i}^{\prime})-\mathcal{G}_{i}(m_{i}^{\prime}+c_{i})).$
Combining the above observations with a standard Taylor series expansion, there exists reals $\{\xi_{i,j}\}_{1\leq i,j\leq n}$, such that:
@@
Recall that all the measures $\{\mu_{i}\}_{1\leq i\leq n}$ are supported on $[-1,1]$, and $g(\cdot)$ takes values in $[-1,1]$ as well. Define a matrix $\mathbf{C}_{n}$ such that $\mathbf{C}_{n}(i,j):=\mathbf{A}_{n}(i,j)\mathcal{G}_{i}^{\prime}(m_{i}+c_{i})$. Then, by our assumption that $\limsup_{n\to\infty}\lVert\mathbf{A}_{n}\rVert<\infty$, we have:
@@
The first and second inequalities in the above sequence follow directly from (5.3). By defining $\mathbf{B}_{n}(i,j):=\mathbf{A}_{n}^{2}(i,j)$, the final inequality follows from
$\lVert\mathbf{B}_{n}\rVert\leq\max_{i=1}^{n}\sum_{\mathnormal{j}=1}^{\mathnormal{n}}\mathbf{B}_{n}(i,j)=\alpha_{n}$ (5.4)
@@
[…]
The final inequality uses standard upper bounds on the Gamma function, see e.g., [30, Theorem 1.1]. We note here that the hidden constants do not depend on $r$. This completes the proof.
@@
$\quad\quad+\big{(}g(\sigma_{j})-2b^{(g)}_{j}\big{)}\big{(}b^{(g)}_{j}-b^{(g),i}_{j}\big{)}+\big{(}b^{(g)}_{j}-b^{(g),i}_{j}\big{)}^{2}\Big{]}.$
@@
$\lesssim\sum_{\mathnormal{i}=1}^{\mathnormal{n}}\gamma_{i}^{2}+\operatorname{\mathbb{E}}\bigg{|}\sum_{i\neq j}\gamma_{i}\gamma_{j}(g(\sigma_{i})-b^{(g)}_{i})b^{(g)}_{i}\Big{[}\big{(}g(\sigma_{j})-2b^{(g)}_{j}\big{)}\big{(}b^{(g)}_{j}-b^{(g),i}_{j}\big{)}+\big{(}b^{(g)}_{j}-b^{(g),i}_{j}\big{)}^{2}\Big{]}\bigg{|}.$
By a similar Taylor expansion as in (5), we note that:
$b^{(g),i}_{j}-b^{(g)}_{j}=-\mathbf{A}_{n}(i,j)\sigma_{i}\mathcal{G}_{j}^{\prime}(m_{j}+c_{j})-\frac{1}{2}\mathbf{A}_{n}(i,j)^{2}\sigma_{i}^{2}\mathcal{G}_{j}^{\prime\prime}(\xi_{i,j}+c_{j})$
for some $\xi_{i,j}\in\mathbb{R}$. As a result, we get
@@
In the last display, we once again use the fact that $\lVert\mathbf{B}_{n}\rVert\leq\max_{i=1}^{n}\sum_{\mathnormal{j}=1}^{\mathnormal{n}}\mathbf{B}_{n}(i,j)=\alpha_{n}$ (see (1.4)). ∎
@@
Let $\|\cdot\|_{q}$ be the $q\to q$ operator norm, with $q\in[2,\infty]$. Also, let $\mathbf{C}_{n}$ be a $p\times p$ matrix. Suppose $\|\mathbf{C}_{n}\|_{q}\leq\theta<1$. Then, $(\mathbf{I}_{n}-\mathbf{C}_{n})^{-1}$ exists and satisfies $\|(\mathbf{I}_{n}-\mathbf{C}_{n})^{-1}\|_{q}\leq\big{(}1-\theta\big{)}^{-1}$.
@@
Suppose $\limsup_{n\to\infty}\lVert\mathbf{A}_{n}\rVert_{q}<1$, for some $q\in[2,\infty]$. Let
@@
Hence, $\mathbf{m}-\mathbf{s}=\mathbf{C}_{n}(\mathbf{m}-\mathbf{s})+\bm{\mathfrak{f}}$, where $\mathbf{C}_{n}(i,j):=\mathbf{A}_{n}(i,j)\psi_{j}^{\prime\prime}(\xi_{j}+c_{j})$ and $\bm{\mathfrak{f}}:=(\mathfrak{f}_{1},\ldots,\mathfrak{f}_{n})$. Note
$\limsup\limits_{n\to\infty}\,\|\mathbf{C}_{n}\|_{q}\leq\limsup\limits_{n\to\infty}\,\lVert\mathbf{A}_{n}\rVert_{q}<1.$
Then, Lemma 5.1 implies $\limsup\limits_{n\to\infty}\|(\mathbf{I}_{n}-\mathbf{C}_{n})^{-1}\|_{q}<\infty$. Thus,
@@
Proof of (c). By using Lemma 2.1 and Lemma 5.1, for any $r\in[2,\infty)$, we have
$\limsup_{n\to\infty}\|(\mathbf{I}_{n}-\mathbf{C}_{n})^{-1}\|_{r}\leq\limsup_{n\to\infty}\|(\mathbf{I}_{n}-\mathbf{C}_{n})^{-1}\|_{\infty}\leq\frac{1}{1-\rho}.$
For any given $r\in[2,\infty)$, by taking $q=r$ in (5.6), we therefore have

### v2

@@ 3 Proof of main results
Proof of 2.10.
The result follows on invoking parts (a) and (b) of 2.6. We show below that the conditions of 2.6 hold, and the error terms converge to 0.
  • MHT condition (2.2)
    Recall from Lemma 2.1 that $\|\mathbf{A}_{n}\|_{4}\leq\|\mathbf{A}\|_{\infty}$, and so it suffices to show the stronger SHT condition (2.3). To this effect, using Bernstein’s inequality for bounded random variables (e.g. Theorem 2.8.4 in [46]) alongside the fact that
    $\operatorname{\mathbb{E}}[\sum_{\mathnormal{j}=1}^{\mathnormal{n}}\mathbf{G}_{n}(i,j)\mid\mathbf{U}]=\sum_{j=1}^{n}\frac{f(U_{i},U_{j})}{n^{\gamma}}\leq n^{1-\gamma},\quad\operatorname{Var}[\sum_{\mathnormal{j}=1}^{\mathnormal{n}}\mathbf{G}_{n}(i,j)\mid\mathbf{U}]\leq n^{1-\gamma},$
    for any $\delta>0$ we have
    $\operatorname{\mathbb{P}}(\|\mathbf{G}_{n}\|_{\infty}>(1+\delta)n^{1-\gamma})$ $=\operatorname{\mathbb{P}}\Big(\max_{\mathnormal{i}=1}^{\mathnormal{n}}\sum_{\mathnormal{j}=1}^{\mathnormal{n}}\mathbf{G}_{n}(i,j)>(1+\delta)n^{1-\gamma}\Big)$
    $\leq 2n\exp\left[-\frac{\delta^{2}}{2(1+\delta/3)}n^{1-\gamma}\right]=o(1).$ (4.9)
    Consequently, we have
    $\|\mathbf{A}_{n}\|_{\infty}=\frac{|\theta|}{n^{1-\gamma}}\|\mathbf{G}_{n}\|_{\infty}\leq|\theta|,$
    where the last inequality holds with high probability using (4). Thus we have verified the SHT condition (2.3) if $|\theta|<1$.
  • SMF condition (2.9)
    Using the fact that $\mathbf{G}_{n}(i,j)\in\{0,1\}$, we have
    $\operatorname{\alpha_{\mathnormal{n}}}=\max_{\mathnormal{i}=1}^{\mathnormal{n}}\sum_{\mathnormal{j}=1}^{\mathnormal{n}}\mathbf{A}_{n}(i,j)^{2}=\max_{\mathnormal{i}=1}^{\mathnormal{n}}\frac{\theta^{2}}{n^{2-2\gamma}}\sum_{\mathnormal{j}=1}^{\mathnormal{n}}\mathbf{G}_{n}(i,j)\leq\frac{\theta^{2}(1+\delta)}{n^{1-\gamma}},$
    where the last step holds with high probability using (4). Since $\gamma<1/2$, we have shown that $\sqrt{n}\alpha_{n}\stackrel{{\scriptstyle P}}{{\to}}0$, and so the SMF condition (2.9) holds.
  • Next, we check each part of (2.18), where we take $q_{i}:=\frac{Q(U_{i})}{\sqrt{n}}$.
    • $\sum_{\mathnormal{i}=1}^{\mathnormal{n}}q_{i}^{2}\to 1$
      This follows as $\sum_{\mathnormal{i}=1}^{\mathnormal{n}}q_{i}^{2}=\frac{1}{n}\sum_{\mathnormal{i}=1}^{\mathnormal{n}}Q(U_{i})^{2}\operatorname*{\xrightarrow{\mathnormal{P}}}\operatorname{\mathbb{E}}Q(U_{1})^{2}=1.$ While (2.18) requires the equality $\sum_{\mathnormal{i}=1}^{\mathnormal{n}}q_{i}^{2}=1$, this can be achieved by re-normalizing $\mathbf{q}$ without changing the limiting variance.
    • $\operatorname{\mathbb{E}}\|\mathbf{A}_{n}\mathbf{q}-\tilde{\lambda}\mathbf{q}\|^{2}=O\Big(\frac{1}{n^{1-\gamma}}\Big)$ for $\tilde{\lambda}:=\theta\lambda$
      A direct expansion gives
      $\operatorname{\mathbb{E}}\|\mathbf{A}_{n}\mathbf{q}-\tilde{\lambda}\mathbf{q}\|^{2}=\operatorname{\mathbb{E}}\Big[\sum_{\mathnormal{i}=1}^{\mathnormal{n}}(\sum_{\mathnormal{j}=1}^{\mathnormal{n}}\mathbf{A}_{n}(i,j)q_{j})^{2}\Big]-2\tilde{\lambda}\operatorname{\mathbb{E}}\Big[\sum_{i,j}\mathbf{A}_{n}(i,j)q_{i}q_{j}\Big]+\tilde{\lambda}^{2}\operatorname{\mathbb{E}}\|\mathbf{q}\|^{2}.$
      We simplify each summand. For the third term we have $\tilde{\lambda}^{2}\operatorname{\mathbb{E}}\|\mathbf{q}\|^{2}=\tilde{\lambda}^{2}$. Proceeding to estimate the second term, direct computations using tower property and the assumption that $Q$ is an eigenfunction of $f$ gives
      $\operatorname{\mathbb{E}}\Big[\sum_{i,j}\mathbf{A}_{n}(i,j)q_{i}q_{j}\Big]$ $=\frac{\theta}{n^{2}}\sum_{i\neq j}\operatorname{\mathbb{E}}[f(U_{i},U_{j})Q(U_{i})Q(U_{j})]=\frac{\tilde{\lambda}(n-1)}{n}.$
      Hence, the middle term is $2\tilde{\lambda}^{2}+O\Big(\frac{1}{n}\Big).$ Finally, to compute the first term, write
      $\operatorname{\mathbb{E}}\Big[(\sum_{\mathnormal{j}=1}^{\mathnormal{n}}\mathbf{A}_{n}(i,j)q_{j})^{2}\mid\mathbf{U}\Big]$ $=\Big[\operatorname{\mathbb{E}}[\sum_{\mathnormal{j}=1}^{\mathnormal{n}}\mathbf{A}_{n}(i,j)q_{j}\mid\mathbf{U}]\Big]^{2}+\operatorname{Var}\Big[\sum_{\mathnormal{j}=1}^{\mathnormal{n}}\mathbf{A}_{n}(i,j)q_{j}\mid\mathbf{U}\Big]$
      $=\frac{\theta^{2}}{n^{2}}\Big[\sum_{j\neq i}f(U_{i},U_{j})q_{j}\Big]^{2}+O\Big(\frac{\sum_{\mathnormal{j}=1}^{\mathnormal{n}}q_{j}^{2}}{n^{2-\gamma}}\Big).$
      Taking a conditional expectation given $U_{i}$ on both sides we get
      $\operatorname{\mathbb{E}}\Big[(\sum_{\mathnormal{j}=1}^{\mathnormal{n}}\mathbf{A}_{n}(i,j)q_{j})^{2}\mid U_{i}\Big]$
      $=$ $\frac{\theta^{2}}{n^{2}}\Big[\operatorname{\mathbb{E}}\big[\sum_{j\neq i}f(U_{i},U_{j})q_{j}\mid U_{i}\big]\Big]^{2}+\frac{\theta^{2}}{n^{2}}\operatorname{Var}\Big[\sum_{j\neq i}f(U_{i},U_{j})q_{j}\mid U_{i}\Big]+O\Big(\frac{\sum_{\mathnormal{j}=1}^{\mathnormal{n}}q_{j}^{2}}{n^{2-\gamma}}\Big)$
      $=$ $\frac{\theta^{2}}{n^{2}}\Big[\frac{n-1}{\sqrt{n}}\lambda Q(U_{i})\Big]^{2}+O\Big(\frac{1}{n^{2}}\Big)+O\Big(\frac{\sum_{\mathnormal{j}=1}^{\mathnormal{n}}Q(U_{j})^{2}}{n^{3-\gamma}}\Big).$
      Here, the last line again uses the fact that $Q$ is the eigenfunction of $f$, along with the fact that $q_{j}=n^{-1/2}Q(U_{j})$. Finally, by taking the expectation over $\mathbf{U}$ and summing over $i$, we have
      $\operatorname{\mathbb{E}}\Big[\sum_{\mathnormal{i}=1}^{\mathnormal{n}}(\sum_{\mathnormal{j}=1}^{\mathnormal{n}}\mathbf{A}_{n}(i,j)q_{j})^{2}\Big]=\frac{\theta^{2}\lambda^{2}(n-1)^{2}}{n^{3}}\operatorname{\mathbb{E}}\Big[\sum_{i=1}^{n}Q^{2}(U_{i})\Big]+O\Big(\frac{1}{n^{1-\gamma}}\Big)=\tilde{\lambda}^{2}+O\Big(\frac{1}{n^{1-\gamma}}\Big).$
      Combining the above bounds we have $\operatorname{\mathbb{E}}\|\mathbf{A}_{n}\mathbf{q}-\tilde{\lambda}\mathbf{q}\|^{2}=O\Big(\frac{1}{n^{1-\gamma}}\Big)$, as claimed.
    • $\|\mathbf{q}\|_{\infty}\to 0$
      It follows from the usual law of large numbers that for i.i.d. random variables $Y_{i}$ with finite mean, we have $\frac{\max_{\mathnormal{i}=1}^{\mathnormal{n}}Y_{i}}{n}\xrightarrow{P}0$. Noting that $\operatorname{\mathbb{E}}Q^{2}(U_{i})=1$, this gives
      $\|\mathbf{q}\|_{\infty}^{2}=\frac{1}{n}\max_{1\leq i\leq n}|Q^{2}(U_{i})|\stackrel{{\scriptstyle P}}{{\to}}0.$
    Having verified all its conditions, the CLTs now follow from 2.6. The additional condition in part (b) of 2.6 also follows from the above bounds, as
    $n\operatorname{\alpha_{\mathnormal{n}}}\|\mathbf{A}_{n}\mathbf{q}-\tilde{\lambda}\mathbf{q}\|^{2}=O_{P}\left(\frac{n}{n^{2(1-\gamma)}}\right)=o_{P}(1),$
    where the last step in the above display uses the fact that $\gamma<\frac{1}{2}$. Note that while we have established all bounds while marginalizing out $\mathbf{U}$, the analogous claims conditional on $\mathbf{U}$ follow from this.
@@
To prove 2.11, we require the following lemmas to check $\sqrt{n}\operatorname{\alpha_{\mathnormal{n}}}=o_{P}(1)$ as well as the MHT condition. Note that unlike the previous examples in the context of graphs, bounding the norm $\|\mathbf{A}_{n}\|_{4}$ by $\|\mathbf{A}_{n}\|_{\infty}$ is suboptimal for spin glasses. We postpone the proofs of these lemmas to Section 6.
@@
Let $\mathbf{Z}$ be a $N\times n$ random matrix whose entries are independent sub-Gaussian random variables with mean zero and variance $1$, and sub-Gaussian norm bounded by a constant. Let $\mathbf{M}_{n}:=\frac{1}{N}\mathbf{Z}^{\top}\mathbf{Z}$, and let $\mathbf{G}_{n}$ be the adjacency matrix of a graph. Suppose $N\geq n$. For $\mathbf{A}_{n}(i,j):=\theta\mathbf{M}_{n}(i,j)\mathbf{G}_{n}(i,j)$ and $r\in[2,\infty]$, the following bound holds:
$\|\mathbf{A}_{n}\|_{r}=O_{P}\Big(\frac{n^{1-1/r}}{\sqrt{N}}\Big).$ (4.10)
While not necessary for our analysis, we also have $\|\mathbf{A}_{n}\|_{r}=O_{P}\Big(\frac{n^{1/r}}{\sqrt{N}}\Big)$ for $r\in[1,2)$.
@@
This result resembles Chevet’s inequality [17, 36], which provides similar $r\to r$ norm bounds for the case when $\mathbf{A}_{n}$ has i.i.d. entries. Note that taking $\mathbf{G}_{n}(i,j)=I(i\neq j)$ in Lemma 4.1 implies that $\mathbf{A}_{n}$ is the off-diagonal part of the sample covariance matrix, and one can easily conclude that
$\|\mathbf{M}_{n}-\mathbf{I}_{n}\|_{r}=O_{P}\Big(\frac{n^{1-1/r}}{\sqrt{N}}\Big)$
for $r\geq 2$ and $N\geq n$. We are not aware of analogous results for covariance matrices, and this lemma may be of independent interest.
@@
Define $\mathbf{A}_{n}$ as in Lemma 4.1. Under the asymptotic scaling $N\gg n^{3/2}$, we have $\max_{\mathnormal{i}=1}^{\mathnormal{n}}\sum_{j=1}^{n}\mathbf{A}_{n}(i,j)^{2}=o_{P}(1/\sqrt{n}).$
Proof of 2.11.
@@
By Lemma 4.2, we have $\sqrt{n}\operatorname{\alpha_{\mathnormal{n}}}=o_{P}(1).$ Also, by Lemma 4.1 with $r=4$, the MHT condition (2.2) holds for all $\theta$. Finally, by Lemma 4.1 with $r=2$, we have $\lVert\mathbf{A}_{n}\rVert=O_{P}\left(\sqrt{\frac{n}{N}}\right),$ and so
$\sqrt{n\alpha_{n}}\lVert\mathbf{A}_{n}\mathbf{q}\rVert=o_{P}(n^{1/4})\lVert\mathbf{A}_{n}\mathbf{q}\rVert\leq o_{P}(n^{1/4})\lVert\mathbf{A}_{n}\rVert=o_{P}(1),$
where the last estimate uses the fact that $N\gg n^{3/2}$. ∎
@@ 5 Proof of main lemmas
Without the loss of generality, assume that the function $g$ is bounded by $1$, i.e. $\|g\|_{\infty}\leq 1$.
Proof of (a). As in the proof of 2.4, we generate $\operatorname{\mathnormal{\bm{\sigma}}}^{\prime}$ as follows: let $I$ be a randomly sampled index from $\{1,2,\ldots,n\}$. Given $I=i$, replace $\sigma_{i}$ with an independently generated $\sigma^{\prime}_{i}$ drawn from the conditional distribution of $\sigma_{i}$ given $\sigma_{j},\ j\neq i$. Recall the definition of $\mathbf{b}^{(g)}$ from 3.1 and let
@@
$f(\operatorname{\mathnormal{\bm{\sigma}}}):=\operatorname{\mathbb{E}}[F(\operatorname{\mathnormal{\bm{\sigma}}},\operatorname{\mathnormal{\bm{\sigma}}}^{\prime})|\operatorname{\mathnormal{\bm{\sigma}}}]=\frac{1}{n}\sum_{\mathnormal{i}=1}^{\mathnormal{n}}\gamma_{i}(g(\sigma_{i})-b_{i}^{(g)}).$
Define $\Delta(\operatorname{\mathnormal{\bm{\sigma}}}):=\operatorname{\mathbb{E}}\left(|f(\operatorname{\mathnormal{\bm{\sigma}}})-f(\operatorname{\mathnormal{\bm{\sigma}}}^{\prime})||F(\operatorname{\mathnormal{\bm{\sigma}}},\operatorname{\mathnormal{\bm{\sigma}}}^{\prime})|\big|\operatorname{\mathnormal{\bm{\sigma}}}\right).$ By [12, Theorem 1.5(ii)], it suffices to show
@@
$f(\operatorname{\mathnormal{\bm{\sigma}}}^{\prime})=\frac{1}{n}\sum_{\mathnormal{i}=1}^{\mathnormal{n}}\gamma_{i}(g(\sigma_{i}^{\prime})-\mathcal{G}_{i}(m_{i}^{\prime}+c_{i})).$
Combining the above observations with a standard Taylor series expansion, there exist reals $\{\xi_{i,j}\}_{1\leq i,j\leq n}$, such that:
@@
Recall that all the measures $\{\mu_{i}\}_{1\leq i\leq n}$ are supported on [1,1][-1,1], and $g(\cdot)$ takes values in [1,1][-1,1] as well. Define a matrix $\mathbf{C}_{n}$ such that $\mathbf{C}_{n}(i,j):=\mathbf{A}_{n}(i,j)\mathcal{G}_{i}^{\prime}(m_{i}+c_{i})$. Then, by our assumption that $\limsup_{n\to\infty}\lVert\mathbf{A}_{n}\rVert<\infty$, we have:
@@
The first and second inequalities in the above sequence follow directly from (5.3). By defining $\mathbf{B}_{n}(i,j):=\mathbf{A}_{n}^{2}(i,j)$, the final inequality follows from using Lemma 2.1 to write
$\lVert\mathbf{B}_{n}\rVert\leq\lVert\mathbf{B}_{n}\rVert_{\infty}=\max_{i=1}^{n}\sum_{\mathnormal{j}=1}^{\mathnormal{n}}\mathbf{B}_{n}(i,j)=\alpha_{n}$ (5.4)
@@
$\leq 2r\int_{0}^{\infty}t^{r-1}e^{-\frac{at^{2}}{\|\bm{\gamma}\|^{2}}}\,dt=\frac{r\lVert\bm{\gamma}\rVert^{r}}{a^{\frac{r}{2}}}\int_{0}^{\infty}t^{\frac{r-2}{2}}e^{-t}\,dt\lesssim\frac{\|\bm{\gamma}\|^{r}}{(2ae)^{\frac{r}{2}}}r^{\frac{r+1}{2}}.$
The final inequality uses standard upper bounds on the Gamma function, see e.g., [40, Theorem 1.1]. We note here that the hidden constants do not depend on $r$. This completes the proof.
@@
$\quad\quad+\big(g(\sigma_{j})-2b^{(g)}_{j}\big)\big(b^{(g)}_{j}-b^{(g),i}_{j}\big)-\big(b^{(g)}_{j}-b^{(g),i}_{j}\big)^{2}\Big].$
@@
$\lesssim\sum_{\mathnormal{i}=1}^{\mathnormal{n}}\gamma_{i}^{2}+\operatorname{\mathbb{E}}\bigg|\sum_{i\neq j}\gamma_{i}\gamma_{j}(g(\sigma_{i})-b^{(g)}_{i})b^{(g)}_{i}\Big[\big(g(\sigma_{j})-2b^{(g)}_{j}\big)\big(b^{(g)}_{j}-b^{(g),i}_{j}\big)|\Big]\bigg|$ (5.5)
$\qquad+\operatorname{\mathbb{E}}\bigg|\sum_{i\neq j}\gamma_{i}\gamma_{j}(g(\sigma_{i})-b^{(g)}_{i})b^{(g)}_{i}\big(b^{(g)}_{j}-b^{(g),i}_{j}\big)^{2}\bigg|.$
By a similar Taylor expansion as in (5), we note that for some $\xi_{i,j}\in\mathbb{R}$:
$b^{(g),i}_{j}-b^{(g)}_{j}=-\mathbf{A}_{n}(i,j)\sigma_{i}\mathcal{G}_{j}^{\prime}(m_{j}+c_{j})+\frac{1}{2}\mathbf{A}_{n}(i,j)^{2}\sigma_{i}^{2}\mathcal{G}_{j}^{\prime\prime}(\xi_{i,j}+c_{j}),\quad|b^{(g),i}_{j}-b^{(g)}_{j}|\lesssim|\mathbf{A}_{n}(i,j)|.$ (5.6)
By using each bound in (5.6) for the second and third summand in (5.5), respectively, we have
@@
$\quad+\frac{1}{2}\operatorname{\mathbb{E}}\bigg|\sum_{i\neq j}\gamma_{i}\gamma_{j}(g(\sigma_{i})-b^{(g)}_{i})b^{(g)}_{i}\big(g(\sigma_{j})-2b^{(g)}_{j}\big)\mathbf{A}_{n}(i,j)^{2}\sigma_{i}^{2}\mathcal{G}_{j}^{\prime\prime}(\xi_{i,j}+c_{j})\bigg|$
[…]
@@
In the last inequality, we bounded (a) the second term by viewing it as a quadratic form, the (b) third term by also viewing it as a quadratic form alongside the operator norm bound of $\mathbf{B}_{n}$ in (5.4). ∎
@@
Let $\|\cdot\|_{q}$ be the $q\to q$ operator norm, with $q\in[2,\infty]$. Also, let $\mathbf{C}_{n}$ be a $n\times n$ matrix. Suppose $\|\mathbf{C}_{n}\|_{q}\leq\theta<1$. Then, $(\mathbf{I}_{n}-\mathbf{C}_{n})^{-1}$ exists and satisfies $\|(\mathbf{I}_{n}-\mathbf{C}_{n})^{-1}\|_{q}\leq\big(1-\theta\big)^{-1}$.
@@
Suppose $\lVert\mathbf{A}_{n}\rVert_{q}<1$, for some $q\in[2,\infty]$. Let
@@
Hence, $\mathbf{m}-\mathbf{s}=\mathbf{C}_{n}(\mathbf{m}-\mathbf{s})+\bm{\mathfrak{f}}$, where $\mathbf{C}_{n}(i,j):=\mathbf{A}_{n}(i,j)\psi_{j}^{\prime\prime}(\xi_{j}+c_{j})$ and $\bm{\mathfrak{f}}:=(\mathfrak{f}_{1},\ldots,\mathfrak{f}_{n})^{\top}$. Note
$\|\mathbf{C}_{n}\|_{q}\leq\lVert\mathbf{A}_{n}\rVert_{q}<1.$
Then, Lemma 5.1 implies $\|(\mathbf{I}_{n}-\mathbf{C}_{n})^{-1}\|_{q}<\infty$. Thus,
@@
Proof of (c). Under the SHT condition (2.3), we have $\|\mathbf{C}_{n}\|_{\infty}\leq\|\mathbf{A}_{n}\|_{\infty}\leq\rho$. By using Lemma 2.1 and Lemma 5.1, for any given $r\in[2,\infty)$, we have
$\|(\mathbf{I}_{n}-\mathbf{C}_{n})^{-1}\|_{r}\leq\|(\mathbf{I}_{n}-\mathbf{C}_{n})^{-1}\|_{\infty}\leq\frac{1}{1-\rho}.$
Then, taking $q=r$ in (5.8) gives